An electric field E can pull an electron out of a Helium atom in the same time that the electron takes
to go around the nucleus once. Then E is
1
Expert's answer
2020-06-22T11:19:00-0400

We can find the required electric field magnitude from the equilibrium of the force of electrostatic attraction and the force caused by the external electric field:


"\frac{1}{4\pi\epsilon_0}\frac{Ze^2}{r^2}=eE,\\\space\\
E=\frac{Ze}{4\pi\epsilon_0r^2}."

For helium:

"Z=2\\

r_{1s}=3.02 \cdot10^{-11}\text{ m}."


"E=3.16\cdot10^{12}\text{ V/m}."